# PAMS: The Perseus Arm Molecular Survey – I. Survey description and first results

### v1

@@ Abstract
The external environments surrounding molecular clouds vary widely across galaxies such as the Milky Way, and statistical samples of clouds from surveys are required to understand them. We present the Perseus Arm Molecular Survey (PAMS), a James Clerk Maxwell Telescope (JCMT) survey of 13CO and C18O (=3–2) of several molecular cloud complexes including W5 and NGC 7538 in the outer Perseus spiral arm situated at and , with a total survey area of 6 deg2. The PAMS data have an effective resolution of 17.2 arcsec, and rms sensitivity of K in 0.3 km s-1 channels. We present a first look at the data, and compare the PAMS regions in the Outer Galaxy with Inner Galaxy regions from the CO Heterodyne Inner Milky Way Plane Survey (CHIMPS), incorporating archival 12CO (3–2) data. By comparing the various CO data with maps of H2 column density from Herschel, we find that the CO-to-H2 column density -factors do not vary significantly between Galactocentric radii of 4 and 10 kpc, and present representative values of and . We find that the emission profiles, size–linewidth and mass–radius relationships of 13CO-traced structures are similar between the Inner and Outer Galaxy. Although PAMS sources are more massive than their Inner Galaxy counterparts for a given size scale, the discrepancy can be accounted for by the Galactic gradient in gas-to-dust mass ratio, uncertainties in the -factors, and selection biases. We have made the PAMS data publicly available, complementing other CO surveys targeting different regions of the Galaxy in different isotopologues and transitions.
@@
Star clusters form within regions of giant molecular clouds (GMCs) where self-gravity is able to overcome opposing physical processes such as thermal pressure, turbulence, and magnetic fields. There are many properties within the interstellar medium (ISM) of a galaxy that vary from one location to another, and these environmental differences could reasonably be expected to leave an imprint upon the process of star formation; in the Milky Way, there are several key differences between the Inner and Outer Galaxy, which we define here as the regions inside and outside of the Sun’s orbit with a Galactocentric radius of kpc (Reid et al., 2019), respectively. For example, the molecular-to-atomic gas ratio, drops from a value of in the central molecular zone (CMZ, kpc) at the centre of the Galaxy, to at kpc (Sofue & Nakanishi, 2016). The strength of the interstellar radiation field (Popescu et al., 2017), dust temperature (Marsh et al., 2017), metallicity (Luck & Lambert, 2011), and the ratio of solenoidal to compressive turbulence within molecular clouds (Rani et al., 2022) all decrease with .
The CMZ is a significantly different star-forming environment from the Galactic disc, with a star formation rate (SFR) an order of magnitude lower than would be expected for the same surface density of molecular gas across a galactic disc (e.g. Longmore et al., 2013; Barnes et al., 2017). The low star formation efficiency (SFE) in the CMZ is accompanied by very large line-of-sight velocity dispersions, and a size–linewidth relationship that is steeper than typically found in the Galactic disc (Kauffmann et al., 2017). These results are generally interpreted as signatures of elevated levels of turbulence that suppress the SFR. Federrath et al. (2016) found that turbulent motions within the CMZ cloud G0.2530.016 are consistent with being dominated by solenoidal modes, which are expected to inhibit star formation (and, conversely, turbulence that is dominated by compressive modes are expected to promote star formation). The authors speculate that shear is responsible for the strong solenoidal modes in the CMZ, caused by the differential rotation of gas that is stronger towards the centres of galaxies.
@@
The Outer Galaxy presents a different star-forming environment to the Inner Galaxy and the CMZ. In addition to the various gradients mentioned above that produce different conditions in the Outer Galaxy, dynamical features also differ in important ways. The interval between the passage of spiral arms is relatively long, and the spiral arms themselves are wider (Reid et al., 2019). The spiral structure itself is less regular, with large deviations from logarithmic-spirals. The corotation radius is located at 7 kpc, outside of which the spiral arm pattern speed exceeds the circular velocity of the gas and stars, and the outer Lindblad resonance is located at 11 kpc (Clarke & Gerhard, 2022). These dynamical effects are good candidates for having an impact upon the star formation process in the Outer Galaxy.
@@
However, the relative sparsity of sources in the Outer Galaxy (i.e. the number of molecular clouds per unit angular area) mean that it is more difficult to justify the expenditure of time on unbiased (i.e. blind) mapping in molecules rarer than 12CO at moderate resolution. Survey data do exist in the Outer Galaxy: the Exeter-FCRAO CO Galactic Plane Survey covers in 12CO and 13CO (1–0) with varying coverage in Galactic latitude at 46 arcseconds resolution, and the Milky Way Imaging Scroll Painting survey (MWISP; Su et al., 2019) covers this region within its staggering and footprint in 12CO, 13CO, and C18O (1–0) at 50 arcseconds resolution. The latest generation of surveys at arcseconds resolution are only now catching up with their Inner Galaxy counterparts: the FOREST unbiased Galactic plane imaging survey with the Nobeyama 45 m telescope (FUGIN; Umemoto et al., 2017), covers 12CO 13CO and C18O (1–0) at 20 arcsecond resolution covers in the third quadrant, which is now also partially () covered by the CHIMPS2 survey (Eden et al. 2020). The CO Large Outer-Galaxy Survey (CLOGS; Eden et al. in preparation) is now extending an area adjoining the CHIMPS2 Outer Galaxy survey in 12CO and 13CO (3–2), and the Outer Galaxy High-Resolution Survey (OGHReS; Urquhart et al., 2024) is also mapping a large area of the third quadrant in the 12CO and 13CO (2–1) with a similar specification as SEDIGISM. In dust continuum the SCUBA-2 Ambitious Sky Survey (SASSy; Nettke et al. 2017, Thompson et al. in preparation) and SASSy-Perseus (Thompson et al. in preparation) surveys cover and , respectively, and have been able to map a very large area in 850 m with a sensitivity comparable to that of ATLASGAL.
The second Galactic quadrant () is relatively unexplored at high angular resolution. In this paper we present the 17 arcsecond-resolution Perseus Arm Molecular Survey (PAMS), a survey of several Outer Galaxy star forming regions in the second quadrant in the 3–2 rotational transition of 13CO and C18O, and incorporating archival 12CO data that partially cover the PAMS regions. The observations are highly complementary to other CO surveys such as CHIMPS, and greatly bolster the available statistics of Outer Galaxy star forming regions. The surveyed regions include some of the most famous star-forming regions within the Perseus arm in the Outer Galaxy, such as W5 and NGC 7538, which lie at kpc.
In Section 2 we describe the observations, and we present the data in Section 3. In Section 4 we compare analyses of the CO-to-molecular hydrogen column density conversion factor, and basic molecular cloud scaling relationships between the Outer Galaxy PAMS data and representative Inner Galaxy data from the CHIMPS survey. We discuss our findings and conclude in Section LABEL:sec:conc.
@@ 2 Observations & Data Reduction
| Region | Area | rms() | |||||||||
| 13CO | C18O | ||||||||||
| 0.3 km s-1 | 0.5 km s-1 | 0.3 km s-1 | 0.5 km s-1 | ||||||||
| (∘) | (∘) | (km s-1) | (kpc) | (kpc) | (kpc) | (deg2) | () | () | () | () | |
| S147/152 | 108.8 | -1.0 | -50.2 | 2.67 | 0.28 | 9.36 | 0.80 | 0.81 | 0.63 | 1.03 | 0.83 |
| G110.5-0.5 | 110.4 | -0.3 | -51.6 | 2.73 | 0.21 | 9.45 | 1.57 | 0.67 | 0.52 | 0.86 | 0.7 |
| S157 | 111.3 | -0.8 | -42.8 | 3.39 | 0.16 | 9.90 | 0.26 | 0.97 | 0.75 | 1.3 | 1.0 |
| NGC 7538 | 111.6 | 0.7 | -52.2 | 2.69 | 0.13 | 9.48 | 0.99 | 0.68 | 0.53 | 0.89 | 0.73 |
| G134.9+1.4 | 134.8 | 1.4 | -40.8 | 1.96 | 0.04 | 9.63 | 0.06 | 0.90 | 0.70 | 1.15 | 0.97 |
| G135.2+1.1 | 135.2 | 1.1 | -44.6 | 1.96 | 0.04 | 9.64 | 0.14 | 0.81 | 0.63 | 1.05 | 0.79 |
| IRAS02327 | 135.5 | 0.3 | -42.8 | 1.94 | 0.04 | 9.63 | 0.03 | 0.54 | 0.42 | 0.72 | 0.56 |
| W5 | 137.8 | 1.3 | -38.4 | 1.96 | 0.04 | 9.69 | 2.42 | 0.66 | 0.51 | 0.83 | 0.69 |
| Total | 6.27 | 0.69 | 0.54 | 0.89 | 0.73 | ||||||
2.1 Observations
@@
We conducted simultaneous basket-woven raster mapping observations of 13CO and C18O (=3–2) at 330.588 GHz and 329.331 GHz using the Heterodyne Array Receiver Program and Auto-Correlation Spectral Imaging System (HARP/ACSIS; Buckle et al., 2009) at the 15-m James Clerk Maxwell Telescope (JCMT) on Mauna Kea, Hawaii. The observations, taken in 2009–10, consist of a series of tiles up to 10201020 arcsec in size of various star-forming regions in the Perseus spiral arm as part of projects M09BU04 and M10BU08. The region list includes well-known complexes such as W5 and NGC 7538, as well as a number of smaller clouds in the region. The targets are listed in Table 1, and the field centres are illustrated upon Planck 857 GHz continuum maps (Planck Collaboration 2020) in Fig. 1. For each tile, these observations took the form of two sets of position-switched scans at right-angles to each other with a quarter array (29.1 arcsec) shift between each scan in a given direction. The same reference positions were used for each tile in a given region, which were checked to ensure they were free of contamination. The 250-MHz bandwidth correlator setting was used with 4096 channels, resulting in a native spectral resolution of 0.06 km s-1. The angular resolution of the JCMT at the observed frequencies is 15 arcsec.
@@
Following standard practice at JCMT, pointing was checked between observations, for which the uncertainty is estimated to be 2 arcseconds in both azimuth and elevation, resulting in a 3 arcsec radial uncertainty. Calibration was performed using the three-load chopper-wheel method (Kutner & Ulich, 1981) during the observations, with which the intensity of the spectra are placed on the corrected antenna temperature () scale. Spectral standards are also monitored throughout observations, and peak and integrated flux densities are generally found to be accurate to within 10 per cent.
@@ 2.2 Data Reduction
Data reduction was performed using orac-dr (Jenness et al., 2015), which is built on the Starlink (Currie et al., 2014) packages kappa (Currie & Berry, 2013), cupid (Berry et al., 2007), and smurf (Chapin et al., 2013), with which we used the REDUCE_SCIENCE_NARROWLINE recipe. For the primary PAMS data products, we give the template reduction parameters in Appendix A. The data were regridded onto the 6 arcsec pixels using an 8-arcsec FWHM Gaussian smooth, meaning that the reduced data cubes have an effective angular resolution of 17.2 arcsec. First order polynomials were used for baseline subtraction. The final cubes were regridded onto 0.3 km s-1-wide velocity channels as our primary data products, but a second version was also produced with 0.5 km s-1 channels to increase compatibility with CHIMPS (Rigby et al., 2016) and CHIMPS2 (Eden et al., 2020). Mosaics of each of the regions were produced using kappa:wcsmosaic with inverse-variance weighting and the sincsinc interpolation kernel. The cubes were astrometrically matched such that the C18O mosaics share the same pixel grid and size as their 13CO counterparts. The individual reduced data cubes are on the corrected antenna temperature () scale, and we converted the larger mosaics to main beam brightness temperature scale by dividing by the main beam efficiency, , where =0.72 at 330 GHz (Buckle et al., 2009).

### v3

@@ Abstract
The external environments surrounding molecular clouds vary widely across galaxies such as the Milky Way, and statistical samples of clouds are required to understand them. We present the Perseus Arm Molecular Survey (PAMS), a James Clerk Maxwell Telescope (JCMT) survey combining new and archival data of molecular-cloud complexes in the outer Perseus spiral arm in 12CO, 13CO, and C18O (=3–2). With a survey area of 8 deg2, PAMS covers well-known complexes such as W3, W5, and NGC 7538 with two fields at and . PAMS has an effective resolution of 17 arcsec, and rms sensitivity of –1.0 K in 0.3 km s-1 channels. Here we present a first look at the data, and compare the PAMS regions in the Outer Galaxy with Inner Galaxy regions from the CO Heterodyne Inner Milky Way Plane Survey (CHIMPS). By comparing the various CO data with maps of H2 column density from Herschel, we calculate representative values for the CO-to-H2 column-density -factors, which are and cm-2 (K km s-1)-1 with a factor of 1.5 uncertainty. We find that the emission profiles, size–linewidth, and mass–radius relationships of 13CO-traced structures are similar between the Inner and Outer Galaxy. Although PAMS sources are slightly more massive than their Inner Galaxy counterparts for a given size scale, the discrepancy can be accounted for by the Galactic gradient in gas-to-dust mass ratio, uncertainties in the -factors, and selection biases. We have made the PAMS data publicly available, complementing other CO surveys targeting different regions of the Galaxy in different isotopologues and transitions.
@@
Star clusters form within regions of giant molecular clouds (GMCs) where self-gravity is able to overcome opposing physical processes such as thermal pressure, turbulence, and magnetic fields. There are many properties within the interstellar medium (ISM) of a galaxy that vary from one location to another, and these environmental differences could reasonably be expected to leave an imprint upon the process of star formation. For example, in the Milky Way, there are several key differences between the Inner and Outer Galaxy, which we define here as the regions inside and outside of the Sun’s orbit with a Galactocentric radius of kpc (Reid et al., 2019), respectively. The molecular-to-atomic gas ratio, drops from a value of in the central molecular zone (CMZ, kpc) at the centre of the Galaxy, to at kpc (Sofue & Nakanishi, 2016). The strength of the interstellar radiation field (Popescu et al., 2017), dust temperature (Marsh et al., 2017), metallicity (Luck & Lambert, 2011), and the ratio of solenoidal to compressive turbulence within molecular clouds (Rani et al., 2022) all decrease with .
The CMZ is a significantly different star-forming environment from the Galactic disc, with a star-formation rate (SFR) an order of magnitude lower than would be expected for the same surface density of molecular gas across a galactic disc (e.g. Longmore et al., 2013; Barnes et al., 2017). The low star-formation efficiency (SFE) in the CMZ is accompanied by very large line-of-sight velocity dispersions, and a size–linewidth relationship that is steeper than typically found in the Galactic disc (Kauffmann et al., 2017). These results are generally interpreted as signatures of elevated levels of turbulence that suppress the SFR. Federrath et al. (2016) found that turbulent motions within the CMZ cloud G0.2530.016 are consistent with being dominated by solenoidal modes, which are expected to inhibit star formation (and, conversely, turbulence that is dominated by compressive modes are expected to promote star formation). The authors speculate that shear is responsible for the strong solenoidal modes in the CMZ, caused by the differential rotation of gas that is stronger towards the centres of galaxies.
@@
The Outer Galaxy presents a different star-forming environment from the Inner Galaxy and the CMZ. In addition to the various gradients mentioned above that produce different conditions in the Outer Galaxy, dynamical features also differ in important ways. The interval between the passage of spiral arms is relatively long, and the spiral arms themselves are wider (Reid et al., 2019). The spiral structure itself is less regular, with large deviations from logarithmic-spirals. The corotation radius is located at 7 kpc, outside of which the spiral arm pattern speed exceeds the circular velocity of the gas and stars, and the outer Lindblad resonance is located at 11 kpc (Clarke & Gerhard, 2022). Although spiral arms do not appear to substantially alter star formation in the Inner Galaxy (e.g. Moore et al., 2012; Eden et al., 2013, 2015; Urquhart et al., 2021; Colombo et al., 2022; Querejeta et al., 2024), these dynamical effects may still have an impact (perhaps indirectly) upon the star-formation process in the Outer Galaxy, where the surface density of molecular clouds and their complexes is much lower.
@@
However, the relative sparsity of sources in the Outer Galaxy (i.e. the number of molecular clouds per unit angular area) mean that it is more difficult to justify the expenditure of time on unbiased (i.e. blind) mapping in molecules rarer than 12CO at moderate resolution. Survey data do exist in the Outer Galaxy: the FCRAO and Exeter-FCRAO CO Galactic Plane Surveys cover in 12CO and 13CO (1–0) with varying coverage in Galactic latitude at 46-arcsecond resolution (Heyer et al., 1998; Wienen et al., 2022, Brunt et al. in prep.); the Forgotten Quadrant survey covers longitudes of with (Benedettini et al., 2020); and the Milky Way Imaging Scroll Painting survey (MWISP; Su et al., 2019) covers this region within its staggering and footprint in 12CO, 13CO, and C18O (1–0) at 50-arcsecond resolution. The latest generation of surveys at -arcsecond resolution is only now catching up with their Inner Galaxy counterparts: the FOREST unbiased Galactic Plane imaging survey with the Nobeyama 45-m telescope (FUGIN; Umemoto et al., 2017), covers 12CO, 13CO, and C18O (1–0) at 20-arcsecond resolution covers in the third quadrant, which is now also partially () covered by the CHIMPS2 survey (Eden et al. 2020). The CO Large Outer-Galaxy Survey (CLOGS; Eden et al. in preparation) is now extending an area adjoining the CHIMPS2 Outer Galaxy survey, primarily in 12CO, but with follow-ups of bright sources in 13CO and C18O(3–2); and the Outer Galaxy High-Resolution Survey (OGHReS; Urquhart et al., 2024) is also mapping a large area of the third quadrant in the 12CO and 13CO (2–1) with a similar specification as SEDIGISM. In dust continuum, the SCUBA-2 Ambitious Sky Survey (SASSy; Nettke et al. 2017, Thompson et al. in preparation) and SASSy-Perseus (Thompson et al. in preparation) surveys cover and , respectively, and have been able to map a very large area in 850 m with a sensitivity comparable to that of ATLASGAL.
The second Galactic quadrant () is relatively unexplored at high angular resolution. In this paper we present the 17-arcsecond-resolution Perseus Arm Molecular Survey (PAMS), a survey of several Outer Galaxy star-forming regions in the second quadrant in the 3–2 rotational transition of 13CO and C18O, and incorporating new and archival 12CO data that cover most of the PAMS regions. The observations are highly complementary to other CO surveys such as CHIMPS/2, COHRS, and CLOGS, and greatly bolster the available statistics of Outer Galaxy star-forming regions. The surveyed regions include some of the most famous star-forming regions within the Perseus arm in the Outer Galaxy, such as W3, W5, and NGC 7538, which lie at kpc.
In Section 2 we describe the observations, and we present the data in Section 3. In Section 4 we compare analyses of the CO-to-molecular-hydrogen column-density conversion factor, and basic molecular-cloud scaling relationships between the Outer Galaxy PAMS data and representative Inner Galaxy data from the CHIMPS survey. We discuss our findings and conclusions in Section 6.
@@ 2 Observations & Data Reduction
| Region | Area | rms() | |||||||||||
| 12CO | 13CO | C18O | |||||||||||
| 0.3 | 0.5 | 0.3 | 0.5 | 0.3 | 0.5 | ||||||||
| km s-1 | km s-1 | km s-1 | km s-1 | km s-1 | km s-1 | ||||||||
| (∘) | (∘) | (km s-1) | (kpc) | (kpc) | (kpc) | (deg2) | () | () | () | () | () | () | |
| W5 | 137.8 | 1.3 | -38.6 | 1.96 | 0.04 | 9.69 | 2.42 | 1.01 | 0.75 | 0.66 | 0.51 | 0.83 | 0.69 |
| IRAS 02327 | 135.5 | 0.3 | -43.1 | 1.96 | 0.04 | 9.65 | 0.03 | – | 0.91 | 0.54 | 0.42 | 0.72 | 0.56 |
| G135.2+1.1 | 135.2 | 1.1 | -44.8 | 1.96 | 0.04 | 9.64 | 0.14 | – | 0.60 | 0.81 | 0.63 | 1.05 | 0.79 |
| G134.9+1.4 | 134.8 | 1.4 | -40.6 | 1.96 | 0.04 | 9.63 | 0.06 | – | – | 0.90 | 0.70 | 1.15 | 0.97 |
| W3 | 133.5 | 0.7 | -45.4 | 1.96 | 0.04 | 9.60 | 1.97 | 0.76 | 0.59 | 0.61 | 0.47 | 0.84 | 0.65 |
| NGC 7538 | 111.6 | 0.7 | -53.1 | 2.69 | 0.13 | 9.48 | 0.99 | – | 0.35 | 0.68 | 0.53 | 0.89 | 0.73 |
| S157 | 111.3 | -0.8 | -43.4 | 3.38 | 0.15 | 9.89 | 0.26 | – | 0.85 | 0.97 | 0.75 | 1.3 | 1.0 |
| G110.5-0.5 | 110.4 | -0.3 | -51.3 | 2.72 | 0.19 | 9.45 | 1.57 | 1.39 | 1.16 | 0.67 | 0.52 | 0.86 | 0.7 |
| S147/152 | 108.8 | -1.0 | -50.8 | 2.81 | 0.22 | 9.44 | 0.80 | 1.00 | 0.84 | 0.81 | 0.63 | 1.03 | 0.83 |
| Total | 8.24 | 1.00 | 0.72 | 0.68 | 0.53 | 0.88 | 0.71 | ||||||
The data presented here as PAMS are comprised of data cubes of the (3–2) transition of 12CO, 13CO, and C18O for nine regions in the Outer Perseus spiral arm. Eight of those regions were originally targeted for PAMS in 13CO and C18O with observing campaigns in 2009–10, and we have supplemented these with archival 12CO data, where available, and obtained new observations of G110 and S152. By incorporating further archival observations of W3, PAMS covers a total of nine regions, with almost complete coverage in the three isotopologues. The observations and data reduction are detailed below.
2.1 13CO and C18O (3–2) observations
@@
We conducted simultaneous basket-woven raster mapping observations of 13CO and C18O (=3–2) at 330.588 GHz and 329.331 GHz using the Heterodyne Array Receiver Program and Auto-Correlation Spectral Imaging System (HARP/ACSIS; Buckle et al., 2009) at the 15-m James Clerk Maxwell Telescope (JCMT) on Mauna Kea, Hawaii. The observations, taken in 2009–10, consist of a series of tiles of up to 10201020 arcsec in size as part of projects M09BU04 and M10BU08. The targets are listed in Table 1, and the field extents are illustrated upon Planck 857 GHz continuum maps (Planck Collaboration 2020) in Fig. 1. For each tile, these observations took the form of two sets of position-switched scans at right angles to each other with a quarter array (29.1 arcsec) shift between each scan in a given direction. The same reference positions were used for each tile in a given region, which were checked to ensure they were free of contamination. The 250-MHz bandwidth correlator setting was used with 4096 channels, resulting in a native spectral resolution of 0.06 km s-1. The native angular resolution of the JCMT at 330 GHz is 15 arcsec.
@@
Following standard practice at JCMT, pointing was checked between observations, for which the uncertainty is estimated to be 2 arcsec in both azimuth and elevation, resulting in a 3 arcsec radial uncertainty. Calibration was performed using the three-load chopper-wheel method (Kutner & Ulich, 1981) during the observations, with which the intensity of the spectra are placed on the corrected antenna temperature () scale. Spectral standards are also monitored throughout observations, and peak and integrated flux densities are generally found to be accurate to within 10 per cent.
2.2 12CO (3–2) observations
New observations of 12CO (3–2) at 345.796 GHz for the G110 and S152 regions were obtained in the summer of 2024 through Programs M23BN003 (PI: Parsons) and S24BP004 (PI: Eden). The setup was similar to the 13CO and C18O observations except for the following differences: the backend (ACSIS) was configured to use 8192 250-MHz-wide channels, all individual tiles were 13201320 arcsec in size, and half-array (58.2 arcsec) spacing was used between rows. The native angular resolution of JCMT at 345 GHz is 14 arcsec.
2.3 Archival observations
We have incorporated further 12CO (=3–2) data from the JCMT archives covering five of the regions: G135.2+1.1, IRAS 02327, S157, NGC 7538, and W5 have almost complete coverage and, while G110 and S152 had partial coverage in the archive, those data were superseded by our new observations. G134.9+1.4 is the only region that has no 12CO (3–2) coverage. The archival data used were M07AU08 (PI: S. Lumsden), M07BH45B, M08BH15, and M09BC12 (PI: J. Williams), M09BN07 (PI: M. Hogerheijde), M09BH09C (PI: J. Di Francesco) and M10BC04 (PI: M. Reid). The NGC 7538 and W5 data were previously published in Fallscheer et al. (2013) and Ginsburg et al. (2011), respectively, while outflows within G110, S152, and S157 formed part of the sample of Maud et al. (2015).
Additional 12CO, 13CO, and C18O (3–2) observations of the W3 complex were incorporated from observing programs M06BU21, M07BH17B, M08BU24 (PI: Moore). The W3 data were originally presented in (Polychroni et al., 2012). Some of these observations date from earlier in the lifetime of the then recently commissioned HARP instrument and have a slightly different observing format, most notably that the scans are not all basket-woven.
@@ 2.4 Data reduction
Data reduction was performed using orac-dr (Jenness et al., 2015), which is built on the Starlink (Currie et al., 2014) packages kappa (Currie & Berry, 2013), cupid (Berry et al., 2007), and smurf (Chapin et al., 2013), with which we used the REDUCE_SCIENCE_NARROWLINE recipe. We give the template reduction parameters in Appendix A. The 13CO and C18O data were regridded onto the 6-arcsec pixels using an 8-arcsec FWHM Gaussian smooth, meaning that the reduced data cubes have an effective angular resolution of 17.2 arcsec. In the vast majority of cases, first-order polynomials were used for baseline subtraction, although fourth-order fits were adopted in some cases. The final cubes were regridded onto 0.3 km s-1-wide velocity channels as our primary data products, but a second version was also produced with 0.5 km s-1 channels to increase compatibility with CHIMPS (Rigby et al., 2016) and CHIMPS2 (Eden et al., 2020).
The 12CO (3–2) data were reduced in an almost identical way using orac-dr, differing only in the use of a 9-arcsec FWHM Gaussian smooth during the regridding, which results in data cubes whose effective angular resolution matches the 17.2-arcsecond resolution of the 13CO and C18O data cubes. Some of the CO data were observed with ACSIS configured with a 1000 MHz bandwidth rather than 250 MHz, which provides a native resolution of 0.42 km s-1, and so the CO data were rebinned to 0.5 km s-1 velocity channels throughout for consistency, with the exception of G110, S152, and part of W3, for which the 0.3 km s-1 cubes are also available.
Mosaics of each of the regions were produced using kappa:wcsmosaic with inverse-variance weighting and the sincsinc interpolation kernel. The cubes were astrometrically matched such that the C18O and 12CO mosaics share the same pixel grid and size as their 13CO counterparts. The individual reduced data cubes are on the corrected antenna temperature () scale, and we converted the larger mosaics to main beam brightness temperature scale by dividing by the main beam efficiency, , where =0.72 at 330 GHz, or at 345 GHz (Buckle et al., 2009). In Fig. 2 we display a combined view111Colour image created using the multicolorfits Python package. of the 12CO (3–2) emission alongside the PAMS 13CO (3–2) emission for NGC 7538, highlighting the utility of the combined data sets. The regions of orange emission trace the highest column densities of CO, while those of yellow emission trace the diffuse envelope of the region.